CORIOLIS-DEPENDENT DIPOLE MOMENT OF THE V_{3} = 2 STATE OF METHANE

\underline{H. SASADA}, K. SUZUMURA, AND C. ISHIBASHI, Department of Physics, Faculty of Science and Technology, Keio University, Yokohama 223, Japan.

Stark effect of the 2ν_{3} band of methane was studied by saturation spectroscopy using an optical cavity absorption cell and a 1.6-μm external cavity semiconductor laser. The magnitude of the determined dipole moments was dependent on the Coriolis sublevels. This dependence agreed with the higher-order dipole moment ^{a},

μ_{α} ∝ ℓ_{β}J_{γ} + J_{β}ℓ_{γ}

where μ is the dipole moment, ℓ are the vibrational angular momenta, J are the total angular momenta, and (α, β, γ) is (x, y, z), (y, z, x), or (z, x, y) components of the molecule-fixed axes.

aM. R. Aliev, V. M. Mikhailov, and J. K. G. Watson, J. Mol. Spectrosc. 118, 544 (1986).